OpenStudy (anonymous):

I need help with understandig this.. please show steps? 4(x - 6) = -2(x - 3)

OpenStudy (anonymous):

open the brackets.

OpenStudy (anonymous):

The first thing you have to do is Distribute the 4 and the -2 can you do that and show me what you get? if you don't know how thats fine i can show you (:

OpenStudy (skullpatrol):

Do you recall the distributive property: $$a(b+c)=ab+ac$$

OpenStudy (uri):

4(x - 6) = -2(x - 3) Solve the brackets 4x-24=-2x+6 Bring all x values together all numericals together 4x+2x=6+24 6x=30 x=30/6 x=5

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):